# Piezoelectric actuation for integrated photonics

### v1

@@ 2.3 Surface and Bulk Acoustic Waves
Besides the quasi-DC (DC-MHz) actuation, when driving at microwave frequencies (GHz), the vibration of the piezoelectric actuator will launch micron to sub-micron wavelength acoustic waves enabling efficient acousto-optic modulation (AOM). AOM has been conventionally adopted in laser systems, spectroscopy, and atomic physics because of its flexibility in controlling the light direction, intensity, and frequency precisely. Depending on the propagation direction of the acoustic waves, they can be mainly classified into surface and bulk acoustic waves.
@@
The traditional AOMs are mostly realized by exciting surface acoustic waves (SAW) through interdigital transducers (IDT) fabricated on the surface of a bulk piezoelectric crystal (e.g., fused silica, LiNbO3, chalcogenide glasses). They mainly work for free-space light with an operating frequency on the order of 100 MHz and consume a substantial electrical power (a few Watts). Thanks to the advances in nano-fabrication and wafer-level deposition of high-quality piezoelectric thin films, it is only within the last decade that the integration of AOM on chip has been successfully demonstrated, extending the modulation frequency beyond 10 GHz [14, 132, 133, 139, 138, 164, 137, 136]. Most of these demonstrations excite SAW by fabricating co-planar IDT electrodes as shown in Fig. 3(a), where an alternative electric field is generated which in turn excites periodically oscillating strain field. Therefore, the wavelength of the acoustic wave is dictated by the period of the IDT. At microwave frequencies, due to the slow acoustic velocity (five orders of magnitude smaller than the speed of light), the acoustic waves have wavelength on the similar order with the optical wavelength, which helps with better mode overlap and interaction strength [14].
@@
Depending on the polarization direction of the propagating wave, different acoustic modes can be identified, such as the Rayleigh wave [14], Lamb wave [165] and Love wave [166], as illustrated in Fig. 3(b-e). While the first two oscillate out of plane, the last one is in plane. In the Rayleigh wave, the particles follow an ellipse trajectory with a long axis in the out of plane direction. It is primarily a shear (transverse) wave but also has a longitudinal component. The Lamb wave is similar as the Rayleigh wave but mainly refers to a plate with finite thickness, such as the released piezoelectric thin film. The Love wave, also known as the Shear Horizontal (SH) wave, can be supported in both released and unreleased configurations. The excitation of these acoustic modes is determined by the alignment of the electric field with the corresponding component in the piezoelectric coefficient tensor, which should be chosen carefully for specific crystal orientations in order for the highest efficiency.
@@
Apart from SAW, by placing electrodes at the top and bottom of the piezoelectric layer, the acoustic waves can be excited vertically towards the substrate, which is the so called High-overtone Bulk Acoustic Waves Resonances (HBAR). As shown in Fig. 4, an acoustic Fabry-Pérot cavity is naturally formed by the top and bottom surfaces of the substrate, which confines tightly the acoustic energy and forms a rich family of acoustic resonant modes. As illustrated in the simulation of Fig. 4(b), an oscillating mechanical displacement wave is excited and occupies uniformly the entire substrate. Depending on the piezoelectric material and the placement of the electrodes, the excitation of both longitudinal [12] and transverse (shear) [167] acoustic waves has been demonstrated. There is no restriction on the thickness of the substrate, which can be either a thick unreleased substrate (e.g., 200-500 m) or a suspended thin film (which is also referred as the Film Bulk Acoustic Waves Resonator, FBAR). HBAR has been ubiquitously deployed in modern wireless systems for high-quality RF filters [168, 169] and oscillators [170, 171], and also coupled with quantum systems such as the superconducting qubits [172, 173, 174] and the Nitrogen Vacancy centers [175, 176]. Most recently, efficient AOM by coupling HBAR with photonic microring resonators has been demonstrated [12, 177].
@@ 2.3.1 Piezo-Optomechanical Model and Coupling Efficiency
The excitation of the SAW and BAW via the piezoelectric actuator can be described either by the Mason model [178, 12, 179] or the Butterworth–Van Dyke (BVD) model [180, 181]. The Mason model describes the acoustic wave based on the transmission line model which captures the physical nature of the acoustic wave propagation and can be used for predicting the electromechanical response based on the structure. Although the Mason model is more general, it is complicated, especially when it is embedded in a complex electric circuit. On the other hand, the BVD model is built on a lumped element circuit, which directly relates with the properties of the specific mechanical mode. It is an empirical model which needs to be fitted with the electromechanical response obtained from either the simulation or the experiment. Nevertheless, the BVD model is one of the most widely used model in the MEMS community.
@@
where (= ) and (= ) are the total loss of the optical and mechanical modes, respectively. A rotating frame around the frequency of the pump light is assumed such that is the relative detune between the pump laser and the optical cavity. () represents the input field in the optical waveguide (microwave transmission line). They are normalized such that () is the photon flux of the pump light (microwave signal) [186]. With the same normalization, () is the average intra-cavity photon number (phonon number).
The above model can be used to describe quantum transduction between optical and mechanical modes by treating and as the annihilation operator. It can also be easily incorporated into other quantum systems, such as the superconducting qubit [183]. On the other hand, the classical AOM can be modeled simply by taking the time average on the field amplitude and removing the hat: , [15, 186]. Moreover, the model is generic to any optomechanical system, regardless of the structure, size, and shape of the optical and mechanical modes, such as the Fabry-Pérot cavity, the microring resonator, or optomechanical crystal. All we need to do is extracting the parameters of the system through simulation and/or experiments. One of the important parameters is the external coupling rate from the mechanical mode to the microwave transmission line with impedance of . Its relation with the electromechanical BVD model has been derived in Refs. [181, 185]:
From the above equation we can see that, larger electromechanical coupling leads to higher external coupling rate.
@@ 3 Ultra-fast Piezoelectric Tuning of Integrated Photonic Devices
In the last decades, integrated photonics has evolved into a mature technology enabling the synthesis, processing and detection of optical signals on-chip. Currently developed photonics integrated circuit (PIC) has enabled nearly all needed functionalities via heterogeneous or hybrid integration of different material platforms [195, 196, 197, 198], and is now utilized at a commercial level for telecommunications and datacenters. A second wave is currently under way in which the nonlinearity of PICs becomes relevant for applications, i.e. integrated nonlinear photonics. The Kerr, Pockels and Brillouin nonlinearities inherent in these waveguides enable new capability for nonlinear optical signal generation and processing [27, 199, 28, 110].
@@
Yet, Si3N4 is all passive, however few active function can still be realized on Si3N4 in combination of other materials, thanks to the maturity and versatility of integrated photonics. For example, modulators can be realized on Si3N4 with integrated 2D-materials (such as graphene) [200, 201, 202, 203], ferroelectric lead-zirconate-titanate (PZT)[91, 75, 88, 52], piezoelectric aluminium nitride [89, 12, 54] and electro-optic LiNbO3 [80, 204, 205]. Multiple approaches have been demonstrated to combine Si3N4 photonic circuits with lasers via hybrid [39, 40, 206, 41, 207, 4] or heterogeneous integration [7, 198]. Today, Si3N4 has emerged as a leading material for integrated linear and nonlinear [27, 198] photonics due to its low loss and superior power handling capability [208]. To date, among all integrated platforms [42], optical losses approaching or even below 1 dB/m have only been demonstrated in Si3N4 PIC [209, 210, 211, 5, 22, 23, 24, 212].
@@ 3.1.1 Actuator Design and DC Tuning
As the well-known piezoelectric material, Aluminum Nitride (AlN) has been widely deployed in modern acoustic radio-frequency (RF) filters due to its mature foundry-level process. We demonstrated for the first time the tuning of a Si3N4 microring resonator by integrating an AlN piezoelectric actuator on top [12, 54]. One typical fabricated device is shown in Fig. 7(a), where a disk shaped actuator overlaps on top of the microring resonator. The bus-waveguide-to-microresonator coupling region is opened in order to prevent any actuation-induced perturbation of light coupling (however, the impact might be very negligible). From the cross-section in Fig. 7(b), 1-m-thick AlN film is sandwiched between between top Aluminum (Al, 100 nm thick) and bottom Molybdenum (Mo, 100 nm thick) metal layers. Molybdenum is chosen to minimize the acoustic impedance with AlN compared to gold or titanium. Ultraow-loss Si3N4 waveguides are fabricated using the photonic Damascene process [213, 214, 24]. The optical waveguide has a dimension of 0.81.8 m2, which is fully buried in a 6-m-thick SiO2 cladding and placed 2 m within the edge of the top metal. The 3-m-thick SiO2 above the waveguide helps to prevent metal absorption from the Mo layer. The entire device sits on a solid 230-m-thick Si substrate, which is called an “unreleased” structure in the MEMS community. The false-colour scanning electron microscopy (SEM) image of the cross-section of the device is shown in Fig. 7(c), including the optical mode in the Si3N4 waveguide.
@@
Polycrystalline AlN thin film and Mo are sputtered on the SiO2 cladding which has been polished via chemical-mechanical planarization (CMP) before the sputtering. The sub-nanometer roughness of the SiO2’s surface helps maintain good crystal orientation of the AlN film, which usually has an X-Ray diffraction (XRD) rocking curve around 2 degree. The AlN has a wurtzite structure with -axis pointing vertically in the positive -axis. Upon applying a positive electrical voltage on the top metal (while bottom metal is grounded), an electric field along the negative -direction is formed. Since AlN has a positive piezoelectric coefficient , it will be squeezed in the -direction and expand horizontally (AlN has a positive Poisson ratio). From the finite-element method (FEM) numerical simulation in Fig. 7(d), extensive (positive) stress and stress gradient are generated around the optical waveguide, and a horizontal mechanical displacement pushes the microring outwards. On the other hand, when applying negative voltages, AlN actuator changes from expanding to shrinking, such that the stress in the above analysis changes sign. In this way, bi-directional tuning can be achieved by reversing the applied voltage’s sign. The optical mode distribution can be similarly simulated, and the fundamental transverse electric (TE) mode is shown in Fig. 7(c). By combining the mechanical and optical simulations, the tuning of the resonant frequency of the optical microring can be estimated by Eq. 12, taking into account the photoelastic and moving boundary effects.
@@
The piezoelectric tuning of the TE mode of the device in Fig. 7 is shown in Fig. 8(a). When applying positive 80 V, the resonant frequency increases by 1.2 GHz, which corresponds to the blue shift. With negative 80 V, the resonance shifts to a lower frequency, showing bi-directional tuning. It can be seen the linewidth of the resonance is maintained during the tuning which means the high optical is unaffected. The estimated intrinsic quality factor, , with integrated AlN actuators is identical to bare microresonators without AlN [215], demonstrating that the monolithically integrated AlN actuators are compatible with ultralow-loss Si3N4 waveguide platform. Moreover, the DC current drawing is maintained below 1 nA, with power consumption at the level of tens of nano-Watts, making it possible for large-scale integration and cryogenic applications. The dependence of resonance shift on the applied voltages is plotted in Fig. 8(b), where a highly linear tuning is obtained up to 140 V with tuning efficiency of 15.7 MHz/V. However, a small hysteresis is observed when sweeping the voltage back and forth. While AlN is non-ferroelectric, the hysteresis can be caused by accumulated charges at the interfaces between AlN and metals and the grain boundaries of the polycrystalline AlN. The tuning range at different wavelengths is shown in Fig. 8(c), which linearly decreases at longer wavelength. This results from a 5.3 MHz change of the microresonator’s FSR (191.0 GHz). The linear dependence of the tuning on wavelength indicates no observable change of the dispersion (where the FSR varies with wavelength). While the capability of controlling the dispersion could find application in versatile programming of the frequency comb’s spectrum [70], the linear tuning of the FSR is preferred in stabilization of the repetition rate of the frequency comb [35].
@@
It can be seen from the stress simulation in Fig. 7(d) that the stress varies dramatically around the edge of the actuator. It is thus necessary to optimize the position of the optical waveguide relative to the actuator. We experimentally fabricated and measured devices with different waveguide positions as shown in Fig. 9. As the waveguide moves from outside of the actuator to inside (negative shift to positive), the tuning efficiency for TE mode increases and saturates as it move more inside, where the stress becomes more uniform. A maximum of 32 MHz/V tuning of TE mode is achieved for +8 m shift. Interestingly, the tuning maintains nearly constant around 20 MHz/V for TM mode due to the anisotropy of the photoelastic effect. Therefore, it is preferred to design the actuator such that the waveguide is well inside the actuator.

### v2

@@ 2.3 Surface and Bulk Acoustic Waves
Besides the quasi-DC (DC-MHz) actuation, when driving at microwave frequencies (GHz), the vibration of the piezoelectric actuator will launch micron to sub-micron wavelength acoustic waves. Depending on the propagation direction of the acoustic waves, they can be classified mainly into surface and bulk acoustic waves. An acoustic resonator functions as a container that confines vibrational waves for either amplification or absorption. The behavior of acoustic waves within such a resonator is governed by the acoustic wave equation and appropriate boundary conditions, which can either be free or fixed. A free boundary condition implies that displacement is at its maximum while stress is zero, typically occurring at the interface with air or vacuum. Conversely, a fixed boundary condition means that displacement is zero and stress is at its peak. At the boundaries between different materials, acoustic waves may be partially or fully reflected due to mismatched acoustic impedance. Acoustic resonance arises when these reflected waves constructively interfere with each other, with the resonator’s cavity length corresponding to an integer multiple of half-wavelengths. The piezoelectric actuator acts as an acoustic source, facilitating the entry of more vibrational waves into the cavity, generally by exerting pressure on one wall or gap within the cavity. [175]. With the acoustic energy enhanced by the acoustic resonances, efficient AOM can be realized. AOM has been conventionally adopted in laser systems, spectroscopy, and atomic physics because of its flexibility in controlling the light direction, intensity, and frequency precisely.
@@
The traditional AOMs are mostly realized by exciting surface acoustic waves (SAW) through IDT fabricated on the surface of a bulk piezoelectric crystal (e.g., fused silica, LiNbO3, chalcogenide glasses). They mainly work for free-space light with an operating frequency on the order of 100 MHz and consume a substantial electrical power (a few Watts). Thanks to the advances in nano-fabrication and wafer-level deposition of high-quality piezoelectric thin films, it is only within the last decade that the integration of AOM on chip has been successfully demonstrated, extending the modulation frequency beyond 10 GHz [14, 141, 142, 148, 147, 176, 146, 145]. Most of these demonstrations excite SAW by fabricating co-planar IDT electrodes as shown in Fig. 3(a), where an alternating electric field is generated which in turn excites periodically oscillating strain field. Therefore, the wavelength of the acoustic wave is dictated by the period of the IDT. At microwave frequencies, due to the slow acoustic velocity (five orders of magnitude smaller than the speed of light), the acoustic waves have wavelength on the similar order with the optical wavelength, which improves mode overlap and interaction strength [14].
@@
Depending on the polarization direction of the propagating soundwave, different acoustic modes can be identified, such as the Rayleigh wave [14], Lamb wave [177] and Love wave [178], as illustrated in Fig. 3(b-e). While the first two oscillate out of plane, the last one is in plane. In the Rayleigh wave, the particles follow an ellipse trajectory with a long axis in the out of plane direction. It is primarily a shear (transverse) wave but also has a longitudinal component. The Lamb wave is similar as the Rayleigh wave but mainly refers to a plate with finite thickness, such as the released piezoelectric thin film. The Love wave, also known as the Shear Horizontal (SH) wave, can be supported in both released and unreleased configurations. The excitation of these acoustic modes is determined by the alignment of the electric field with the corresponding component in the piezoelectric coefficient tensor, which should be chosen carefully for specific crystal orientations in order to achieve the highest efficiency.
@@
Apart from SAW, by placing electrodes at the top and bottom of the piezoelectric layer, the acoustic waves can be directed vertically towards the substrate, exciting so-called High-overtone Bulk Acoustic Waves Resonances (HBAR). As shown in Fig. 4, an acoustic Fabry-Pérot cavity is naturally formed by the top and bottom surfaces of the substrate, which confines tightly the acoustic energy and forms a rich family of acoustic resonant modes. As illustrated in the simulation of Fig. 4(b), an oscillating mechanical displacement wave is excited and occupies uniformly the entire substrate. Depending on the piezoelectric material and the placement of the electrodes, the excitation of both longitudinal [12] and transverse (shear) [179] acoustic waves has been demonstrated. There is no restriction on the thickness of the substrate, which can be either a thick unreleased substrate (e.g., 200-500 m) or a suspended thin film (which is also referred as the Film Bulk Acoustic Waves Resonator, FBAR). HBAR has been ubiquitously deployed in modern wireless systems for high-quality RF filters [180, 181] and oscillators [182, 183], and also coupled with quantum systems such as the superconducting (SC) qubits [184, 185, 186] and the Nitrogen Vacancy centers [187, 188]. Most recently, efficient AOM by coupling HBAR with photonic microring resonators has been demonstrated [12, 189].
@@ 2.3.1 Piezo-Optomechanical Model and Coupling Efficiency
The excitation of the SAW and bulk acoustic waves (BAW) via the piezoelectric actuator can be described either by the Mason model [190, 12, 191] or the Butterworth–Van Dyke (BVD) model [192, 193]. The Mason model describes the acoustic wave based on the transmission line model which captures the physical nature of the acoustic wave propagation and can be used for predicting the electromechanical response based on the structure. Although the Mason model is more general, it is complicated, especially when it is embedded in a complex electric circuit. On the other hand, the BVD model is built on a lumped element circuit, which directly relates with the properties of the specific mechanical mode. It is an empirical model, which needs to be fitted with the electromechanical response obtained from either the simulation or the experiment. Nevertheless, the BVD model is one of the most widely used models in the MEMS community.
@@
where (= ) and (= ) are the total loss of the optical and mechanical modes, respectively. A rotating frame around the frequency of the pump light is adopted such that is the relative detuning between the pump laser and the optical cavity. () represents the input field in the optical waveguide (microwave transmission line). They are normalized such that () is the photon flux of the pump light (microwave signal) [198]. With the same normalization, () is the average intra-cavity photon number (phonon number).
The above model can be used to describe quantum transduction between optical and mechanical modes by treating and as the annihilation operator. It can also be easily incorporated into other quantum systems, such as the SC qubit [195]. On the other hand, the classical AOM can be modeled simply by taking the time average on the field amplitude and removing the hat: , [15, 198]. Moreover, the model is generic to any optomechanical system, regardless of the structure, size, and shape of the optical and mechanical modes, such as the Fabry-Pérot cavity, the microring resonator, or optomechanical crystal. All we need to do is extracting the parameters of the system through simulation and/or experiments. One of the important parameters is the external coupling rate from the mechanical mode to the microwave transmission line with impedance of . Its relation with the electromechanical BVD model has been derived in Refs. [193, 197]:
From the above equation we can see that, larger electromechanical coupling leads to stronger external coupling rate.
@@ 3 Ultra-fast Piezoelectric Tuning of Integrated Photonic Devices
In the last decades, integrated photonics has evolved into a mature technology enabling the synthesis, processing and detection of optical signals on-chip. Currently developed PICs have enabled nearly all core functionalities via heterogeneous or hybrid integration of different material platforms [207, 208, 209, 210], and is now utilized at a commercial level for telecommunications and datacenters. A second wave is currently under way in which the nonlinearity of PICs becomes relevant for applications, i.e. integrated nonlinear photonics. The Kerr, Pockels and Brillouin nonlinearities inherent in these waveguides enable new capability for nonlinear optical signal generation and processing [27, 211, 28, 113].
@@
Yet, Si3N4 is all passive, however few active function can still be realized on Si3N4 in combination of other materials, thanks to the maturity and versatility of integrated photonics. For example, modulators can be realized on Si3N4 with integrated 2D-materials (such as graphene) [212, 213, 214, 215], ferroelectric PZT [93, 77, 90, 52], piezoelectric aluminium nitride [91, 12, 54] and electro-optic LiNbO3 [82, 216, 217]. Multiple approaches have been demonstrated to combine Si3N4 photonic circuits with lasers via hybrid [39, 40, 218, 41, 219, 4] or heterogeneous integration [7, 210]. Today, Si3N4 has emerged as a leading material for integrated linear and nonlinear [27, 210] photonics due to its low loss and superior optical power handling capability [157]. To date, among all integrated platforms [42], optical losses below 1 dB/m have been demonstrated in Si3N4 PIC [220, 221, 222, 5, 22, 23, 24, 223].
@@ 3.1.1 Actuator Design and DC Tuning
As the well-known piezoelectric material, AlN has been widely deployed in modern acoustic RF filters due to its mature foundry-level process. We demonstrated for the first time the tuning of a Si3N4 microring resonator by integrating an AlN piezoelectric actuator on top [12, 54]. One typical fabricated device is shown in Fig. 7(a), where a disk shaped actuator overlaps on top of the microring resonator. The bus-waveguide-to-microresonator coupling region is opened in order to prevent any actuation-induced perturbation of light coupling (however, the impact might be very negligible). From the cross-section in Fig. 7(b), 1-m-thick AlN film is sandwiched between between top Aluminum (Al, 100 nm thick) and bottom Molybdenum (Mo, 100 nm thick) metal layers. Molybdenum is chosen to minimize the acoustic impedance with AlN compared to gold or titanium. Ultraow-loss Si3N4 waveguides are fabricated using the photonic Damascene process [224, 225, 24]. The optical waveguide has a dimension of 0.81.8 m2, which is fully buried in a 6-m-thick SiO2 cladding and placed 2 m within the edge of the top metal. The 3-m-thick SiO2 above the waveguide helps to prevent metal absorption from the Mo layer. The entire device sits on a solid 230-m-thick Si substrate, which is called an “unreleased” structure in the MEMS community. The false-colour SEM image of the cross-section of the device is shown in Fig. 7(c), including the optical mode in the Si3N4 waveguide.
@@
Polycrystalline AlN thin film and Mo are sputtered on the SiO2 cladding which has been polished via chemical-mechanical planarization (CMP) before the sputtering. The sub-nanometer roughness of the SiO2’s surface helps maintain good crystal orientation of the AlN film, which usually has an X-Ray diffraction (XRD) rocking curve peak width around 2 degrees. The AlN has a wurtzite structure with -axis pointing vertically in the positive -axis. Upon applying a positive electrical voltage on the top metal (while bottom metal is grounded), an electric field along the negative -direction is formed. Since AlN has a positive piezoelectric coefficient , it will be squeezed in the -direction and expand horizontally (AlN has a positive Poisson ratio). From the FEM numerical simulation in Fig. 7(d), extensive (positive) stress and stress gradient are generated around the optical waveguide, and a horizontal mechanical displacement pushes the microring outwards. On the other hand, when applying negative voltages, AlN actuator changes from expanding to shrinking, such that the stress in the above analysis changes sign. In this way, bi-directional tuning can be achieved by reversing the applied voltage’s sign. The optical mode distribution can be similarly simulated, and the fundamental transverse electric (TE00) mode is shown in Fig. 7(c). By combining the mechanical and optical simulations, the tuning of the resonant frequency of the optical microring can be estimated by Eq. 13, taking into account the photoelastic and moving boundary effects.
@@
The piezoelectric tuning of the TE00 mode of the device in Fig. 7 is shown in Fig. 8(a). When applying positive 80 V, the resonant frequency increases by 1.2 GHz, which corresponds to the blue shift. With negative 80 V, the resonance shifts to a lower frequency, showing bi-directional tuning. It can be seen the linewidth of the resonance is maintained during the tuning which means the high optical is unaffected. The estimated intrinsic quality factor, , with integrated AlN actuators is identical to bare microresonators without AlN [226], demonstrating that the monolithically integrated AlN actuators are compatible with ultralow-loss Si3N4 waveguide platforms. Moreover, the DC current draw is maintained below 1 nA, with power consumption at the level of tens of nano-Watts, allowing large-scale integration and cryogenic applications. The dependence of resonance shift on the applied voltages is plotted in Fig. 8(b), where a highly linear tuning is obtained up to 140 V with tuning efficiency of 15.7 MHz/V. However, a small hysteresis is observed when sweeping the voltage back and forth. While AlN is non-ferroelectric, the hysteresis can be caused by accumulated charges at the interfaces between AlN and metals and the grain boundaries of the polycrystalline AlN. The tuning range at different wavelengths is shown in Fig. 8(c), which linearly decreases at longer wavelength. This results from a 5.3 MHz change of the microresonator’s FSR (191.0 GHz). The linear dependence of the tuning on wavelength indicates no observable change of the dispersion (where the FSR varies with wavelength). While the capability of controlling the dispersion could find application in versatile programming of the frequency comb’s spectrum [72], the linear tuning of the FSR is preferred for stabilization of the repetition rate of the frequency comb [35].
@@
It can be seen from the stress simulation in Fig. 7(d) that the stress varies dramatically around the edge of the actuator. It is thus necessary to optimize the position of the optical waveguide relative to the actuator. We experimentally fabricated and measured devices with different waveguide positions as shown in Fig. 9. As the waveguide moves from outside of the actuator to inside (negative shift to positive), the tuning efficiency for TE00 mode increases and saturates as it move more inside, where the stress becomes more uniform. A maximum of 32 MHz/V tuning of TE00 mode is achieved for +8 m shift. Differently, the tuning maintains nearly constant around 20 MHz/V for TM00 mode due to the anisotropy of the photoelastic effect. Therefore, it is preferred to place the waveguide well inside the actuator for better tuning efficiency.